### v1

@@ 1 Introduction
Quantization methods can be categorized into two primary categories: Quantization Aware Training (QAT) [35, 15, 45, 46, 55, 20, 52] and Post-Training Quantization (PTQ) [24, 31, 7, 26, 16, 47, 42]. In general terms, QAT is designed to globally minimize the conventional training loss of the model for quantization parameters. It involves tackling a formidable nonconvex minimization problem with a discrete nature. QAT requires an all-encompassing training pipeline and computational cost that is at least on par with regular full-precision models. In contrast, PTQ directly applies low-precision calibration to a pre-trained full-precision model. Computationally, PTQ aims to identify an optimal quantized model locally by minimizing a simplified surrogate loss, so it enjoys significantly reduced algorithmic complexity and appears to be a faster and more resource-efficient process. On the downside, PTQ suffers a heavier performance degradation than QAT, especially when it comes to low-bit quantization of Vision Transformers (ViTs), which has received much recent attention [6, 39, 29, 30].
@@
Contribution. In this work, we introduce COMQ, a post-training quantization method that performs uniform weight quantization on a layer-by-layer basis (see Figure 1). Similarly to existing works [31, 16, 7, 54], our main objective is to minimize the layerwise squared error $\|\mathbf{X}\mathbf{W}_{q}-\mathbf{X}\mathbf{W}\|^{2}$ with respect to quantized weights $\mathbf{W}_{q}$. To efficiently address optimization, COMQ enforces the decomposition $\mathbf{W}_{q}=\delta\cdot\mathbf{Q}$ with $\delta$ being the full-precision scalar(s) and $\mathbf{Q}$ storing the integer bit-codes, and it then minimizes this error over the new variables through a coordinate-wise minimization procedure. Adhering to a greedy selection rule, we select one variable at a time, whether scaling factor or bit-code, to update while maintaining the others at their most recent states. Unlike recent works [7, 24, 31] that require back-propagation or estimation of the Hessian inverse to minimize the reconstruction error, COMQ solves a sequence of minimization of univariate quadratic functions which enjoy closed-form minimizers. This leads to backpropagation-free iterations that primarily involve dot products and rounding operations without using any hyperparameters.
@@
Notation. Throughout this paper, we denote vectors by bold small letters and matrices by bold capital ones. For any two vectors $\mathbf{x},\mathbf{y}\in\mathbb{R}^{n}$, $\langle\mathbf{x},\mathbf{y}\rangle:=\sum_{i=1}^{n}x_{i}y_{i}$ is their inner product. We denote by $\|\mathbf{x}\|:=\sqrt{\langle\mathbf{x},\mathbf{x}\rangle}$ the Euclidean norm. Similarly for two matrices $\mathbf{X},\mathbf{Y}\in\mathbb{R}^{m\times n}$, the inner product is given by $\langle\mathbf{X},\mathbf{Y}\rangle:=\sum_{i=1}^{m}\sum_{j=1}^{n}X_{i,j}Y_{i,j}$, and $\|\mathbf{X}\|:=\sqrt{\langle\mathbf{X},\mathbf{X}\rangle}$ is the Frobenius norm.
@@ 2.1 Post-training quantization
Some initiatives have been undertaken to develop backpropagation-free algorithms for weight quantization. OBC [7] adapts the Optimal Brain Surgeon framework to the setting of layer-wise PTQ. OBC quantizes one weight using an analytical formula at a time and updates all remaining weights after each step. Compared with OBC, our proposed COMQ method is simpler because it only updates one parameter in each step and keeps the others untouched. Additionally, unlike OBC, COMQ does not require inverting the Hessian matrix of the layer-wise squared error in quantizing the weight. Zhang et al. [54] introduced GPFQ to efficiently learn layer-wise bit-codes sequentially, assuming that the floating scaling factors are already learned. In its practical implementation, determining the appropriate scalars requires a trial-and-error process.
@@ 2.2 Coordinate descent
In the PTQ problem, each quantized weight is represented as $\mathbf{W}_{q}=\delta\cdot\mathbf{Q}$. We consider the scale $\delta$ and all elements in $\mathbf{Q}$ as coordinates. In one step, we update one element in $\mathbf{Q}$ or the scale $\delta$ while fixing all other coordinates. Each subproblem in the coordinate descent method is a univariate optimization problem and has a closed-form solution in our settings. The order of solving subproblems affects the optimization performance. Cyclic order is the most widely used, cyclically iterating through the coordinates. However, the minimization problem for quantization is non-convex and exhibits a discrete nature, challenging the optimality of cyclic order. In this work, we propose to update the parameters in a carefully designed order to achieve a reduced quantization loss compared to that of cyclic order [44].
@@ 3 Proposed Method
We consider a linear layer with matrix weight $\mathbf{W}$ and matrix input $\mathbf{X}$. For the weight $\mathbf{W}$, we aim to find the quantized weights $\mathbf{W}_{q}$ that minimize the following function
$\min_{\mathbf{W}_{q}\in\mathcal{W}}\;\|\mathbf{X}\mathbf{W}_{q}-\mathbf{X}\mathbf{W}\|^{2},$ (1)
where $\mathcal{W}$ is an appropriate set of all feasible quantized weights. We propose to solve the problem (1) by the coordinate descent algorithm. We regard all elements in the quantized weight and the scale factor as coordinates. In each step, we only solve a univariate optimization problem with respect to only one coordinate. By iteratively computing the quantized weights and corresponding scale factors, we finally find a proper quantized weight $\mathbf{W}_{q}$ that has a small quantization error and maintains good accuracy.
In this section, we will show our COMQ method for two quantization scenarios: per-layer quantization and per-channel quantization. All quantized weights with in the same layer share a common scale factor for the per-layer quantization, while different columns of the quantized weight use different scale factors for the per-channel quantization.
@@ 3.1 Per-layer Quantization
In this subsection, we present our COMQ method for $b$-bit per-layer quantization, where the whole quantized weight matrix $\mathbf{W}_{q}$ shares the scale factor $\delta$. The per-layer quantization will use the uniform quantization since the whole weight matrix $W$ is nearly centered at zero. Recall that, the $b$-bit uniform quantization takes the bit-code set of $\mathbb{S}=\{-2^{b-1},\ldots,2^{b-1}-1\}$.
The quantized weight $\mathbf{W}_{q}$ can be written as
$\mathbf{W}_{q}=\delta\mathbf{Q}\in\mathbb{R}^{m\times n},$
@@
where $\mathbf{Q}$ is the integer bit-code matrix for $\mathbf{W}_{q}$, and $\mathbf{Q}_{i,j}\in\mathbb{S}=\{-2^{b-1},\ldots,2^{b-1}-1\}$. The optimization problem (1) reduces to
$\min_{\delta\in\mathbb{R},\mathbf{Q}_{i,j}\in\mathbb{S}}\;\|\delta\mathbf{X}\mathbf{Q}-\mathbf{X}\mathbf{W}\|^{2}.$ (2)
@@
It holds that $\mathbf{X}\mathbf{Q}=(\mathbf{X}\mathbf{q}_{1},\ldots,\mathbf{X}\mathbf{q}_{n})$ and $\mathbf{X}\mathbf{W}=(\mathbf{X}\mathbf{w}_{1},\ldots,\mathbf{X}\mathbf{w}_{n})$, where $\mathbf{q}_{i}$ and $\mathbf{w}_{i}$ are the $i$th column of $\mathbf{Q}$ and $\mathbf{W}$ respectively. Therefore, we can rewrite the problem (2) as
$\min_{\delta,\mathbf{q}_{1},\ldots,\mathbf{q}_{n}}\sum_{i=1}^{n}\|\delta\mathbf{X}\mathbf{q}_{i}-\mathbf{X}\mathbf{w}_{i}\|^{2}.$ (3)
@@
After fixing the scale factor $\delta$, the terms $\|\delta\mathbf{X}\mathbf{q}_{i}-\mathbf{X}\mathbf{w}_{i}\|^{2}$ for $i=1,\ldots,n$ are independent of each other. Hence, we can consider the columns $\mathbf{q}_{1},\ldots,\mathbf{q}_{n}$ separately. We apply the coordinate descent method to solve the above problem (2). Specifically, we regard the scale factor $\delta$ and all elements in $\mathbf{Q}$ as coordinates. In each iteration, we sequentially update the coordinates across all columns in the following order:
$\{(\mathbf{q}_{i})_{1}\}_{i=1}^{n}\Rightarrow\cdots\Rightarrow\{(\mathbf{q}_{i})_{m}\}_{i=1}^{n}\Rightarrow\delta.$ (4)
@@
Initialization. At the beginning of our COMQ method, the scale factor $\delta^{0}$ should be properly initialized to capture the range of the weight matrix $\mathbf{W}$. Typically, the maximum value of $|\mathbf{W}|$ is the default selection in most quantization problems. However, this choice overlooks the impact of outliers in the weights on the overall quantization error. In order to smooth out these outliers, we consider the average infinity norm of weights across all columns $\mathbf{W}$. For the $b$ bits uniform quantization, we use the initialization $\delta^{0}=\frac{1}{2^{b-1}}\frac{\sum_{1\leq i\leq n}\|\mathbf{w}_{i}\|_{\infty}}{n}$. Then the matrix $\mathbf{Q}$ is initialized as $\mathbf{Q}^{0}=\frac{\mathbf{W}}{\delta^{0}}$. Note that the initialization $\mathbf{Q}^{0}$ above is not an actual bit-code matrix as $\mathbf{W}$ is float, but it will become feasible after the 1st iteration as will be shown below.
Update. Let $\mathbf{Q}^{k-1},\delta^{k-1}$ be the parameters after $k-1$ iterations. The updated coordinate $(\mathbf{q}_{i}^{k})_{j}$ can be obtained by solving the following problem
$\min_{q\in\mathbb{S}}\|q\delta^{k-1}\mathbf{x}_{j}-\mathbf{s}^{k}_{i,j}\|^{2},$ (5)
@@
$(\mathbf{q}_{i}^{k})_{j}=\text{clip}\left(\left\lfloor\frac{\left<\delta^{k-1}\mathbf{x}_{j},\mathbf{s}^{k}_{i,j}\right>}{\|\delta^{k-1}\mathbf{x}_{j}\|^{2}}\right\rceil,-2^{b-1},2^{b-1}-1\right).$ (6)
We compute the updated coordinates in $\mathbf{Q}$ sequentially using the order in (4). After obtaining the new bit-code matrix $\mathbf{Q}^{k}$, the scale factor can be updated by solving the following problem
$\min_{\delta}\|\delta\mathbf{X}\mathbf{Q}^{k}-\mathbf{X}\mathbf{W}\|^{2}.$
@@
$\delta^{k}=\frac{\left<\mathbf{XQ}^{k},\mathbf{XW}\right>}{\|\mathbf{XQ}^{k}\|^{2}}.$ (7)
@@
Algorithm 1 COMQ for per-layer quantization
0:  A linear layer with weight $\mathbf{W}\in\mathbb{R}^{m\times n}$, input $\mathbf{X}$, and number of iterations $K$.
1:  Initialize $\delta^{0}$ and $\mathbf{Q}^{0}$
@@
3:     for $j=1,\ldots,m$ do
4:        Find updated coordinates $(\mathbf{q}_{i}^{k})_{j}$ for $i=1,\ldots,n$ as in (6)
@@
6:     Find the new scale factor $\delta^{k}$ as in (7)
@@
8:  Let $\mathbf{W}_{q}=\delta^{K}\mathbf{Q}^{K}$
8:  Quantized weight $\mathbf{W}_{q}$.

### v3

@@ 1 Introduction
Quantization methods can be categorized into two primary categories: Quantization Aware Training (QAT) [37, 17, 47, 55, 54, 58, 22, 53] and Post-Training Quantization (PTQ) [26, 33, 9, 28, 18, 48, 44]. In general terms, QAT is designed to globally minimize the conventional training loss of the model for quantization parameters. It involves tackling a formidable nonconvex minimization problem with a discrete nature. QAT requires an all-encompassing training pipeline and computational cost that is at least on par with regular full-precision models. In contrast, PTQ directly applies low-precision calibration to a pre-trained full-precision model. Computationally, PTQ aims to identify an optimal quantized model locally by minimizing a simplified surrogate loss, so it enjoys significantly reduced algorithmic complexity and appears to be a faster and more resource-efficient process. On the downside, PTQ suffers a heavier performance degradation than QAT, especially when it comes to low-bit quantization of Vision Transformers (ViTs), which has received much recent attention [8, 41, 31, 32].
@@
Contribution. In this work, we introduce COMQ, a post-training quantization method that performs uniform weight quantization on a layer-by-layer basis (see Figure 1). Similarly to existing works [33, 18, 9, 57], our main objective is to minimize the layerwise squared error $\|\boldsymbol{X}\boldsymbol{W}_{q}-\boldsymbol{X}\boldsymbol{W}\|^{2}$ with respect to quantized weights $\boldsymbol{W}_{q}$. To efficiently address optimization, COMQ enforces the decomposition $\boldsymbol{W}_{q}=\delta\cdot\boldsymbol{Q}$ with $\delta$ being the full-precision scalar(s) and $\boldsymbol{Q}$ storing the integer bit-codes, and it then minimizes this error over the new variables through a coordinate-wise minimization procedure. Adhering to a greedy selection rule, we select one variable at a time, whether scaling factor or bit-code, to update while maintaining the others at their most recent states. Unlike recent works [9, 26, 33] that require back-propagation or estimation of the Hessian inverse to minimize the reconstruction error, COMQ solves a sequence of minimization of univariate quadratic functions which enjoy closed-form minimizers. This leads to backpropagation-free iterations that primarily involve dot products and rounding operations without using any hyperparameters.
@@
Notations. Throughout this paper, we denote vectors with bold small letters and matrices with bold capital ones. For any two vectors $\boldsymbol{x},\boldsymbol{y}\in\mathbb{R}^{n}$, $\langle\boldsymbol{x},\boldsymbol{y}\rangle:=\boldsymbol{x}^{\top}\boldsymbol{y}=\sum_{i=1}^{n}x_{i}y_{i}$ is their inner product. We denote by $\|\boldsymbol{x}\|:=\sqrt{\langle\boldsymbol{x},\boldsymbol{x}\rangle}$ the Euclidean norm and denote by $\|\boldsymbol{x}\|_{\infty}=\max_{i=1,\ldots,n}|x_{i}|$ the $l_{\infty}$ norm. Similarly for two matrices $\boldsymbol{X},\boldsymbol{Y}\in\mathbb{R}^{m\times n}$, the inner product is given by $\langle\boldsymbol{X},\boldsymbol{Y}\rangle:=\sum_{i=1}^{m}\sum_{j=1}^{n}X_{i,j}Y_{i,j}$, and $\|\boldsymbol{X}\|:=\sqrt{\langle\boldsymbol{X},\boldsymbol{X}\rangle}$ is the Frobenius norm. Moreover, we denote the outer product of two vectors by $\boldsymbol{x}\otimes\boldsymbol{y}:=\boldsymbol{x}\boldsymbol{y}^{\top}\in\mathbb{R}^{n\times n}$. $\boldsymbol{x}\odot\boldsymbol{y}:=(x_{1}y_{1},\dots,x_{n}y_{n})\in\mathbb{R}^{n}$ denotes the Hadamard (or element-wise) product, and $\boldsymbol{x}\oslash\boldsymbol{y}:=(x_{1}/y_{1},\dots,x_{n}/y_{n})\in\mathbb{R}^{n}$ denotes the Hadamard division. Finally, for any positive integer $n$, $[n]:=\{1,\dots,n\}$ denotes the set of integers up to $n$.
@@ 2.1 Post-training quantization
Some initiatives have been undertaken to develop backpropagation-free algorithms for weight quantization. OBC [9], also known as OPTQ [10] for quantizing large language models (LLMs), adapts the Optimal Brain Surgeon framework to the setting of layer-wise PTQ. OBC quantizes one weight using an analytical formula at a time and updates all remaining weights after each step. Compared with OBC, our proposed COMQ method is simpler because it only updates one parameter in each step and keeps the others untouched. Additionally, unlike OBC, COMQ does not require inverting the Hessian matrix of the layer-wise squared error in quantizing the weight. Zhang et al. [57] introduced GPFQ to efficiently learn layer-wise bit-codes sequentially, assuming that the floating scaling factors are already learned. In its practical implementation, determining the appropriate scalars requires a trial-and-error process. In parallel to our research, Behdin et al. explored a cyclic coordinate descent approach, QuantEease [1], for PTQ. While their work focuses on LLMs, our research targets ViTs with a different algorithmic implementation. Specifically, we propose a greedy update order and introduce learnable scaling factors, which we believe offer unique benefits in terms of performance.
@@ 2.2 Coordinate descent
In the PTQ problem, each quantized weight is represented as $\boldsymbol{W}_{q}=\delta\cdot\boldsymbol{Q}$. We consider the scale $\delta$ and all elements in $\boldsymbol{Q}$ as coordinates. In one step, we update one element in $\boldsymbol{Q}$ or the scale $\delta$ while fixing all other coordinates. Each subproblem in the coordinate descent method is a univariate optimization problem and has a closed-form solution in our settings. The order of solving subproblems affects the optimization performance. Cyclic order is the most widely used, cyclically iterating through the coordinates. However, the minimization problem for quantization is non-convex and exhibits a discrete nature, challenging the optimality of cyclic order. In this work, we propose to update the parameters in a carefully designed order to achieve a reduced quantization loss compared to that of cyclic order [46].
@@ 3 Proposed Method
We consider a linear layer with matrix weight $\boldsymbol{W}$ and matrix input $\boldsymbol{X}$. For the weight $\boldsymbol{W}$, we aim to find the quantized weights $\boldsymbol{W}_{q}$ that minimize the following function
$\min_{\boldsymbol{W}_{q}\in\mathcal{W}}\;\|\boldsymbol{X}\boldsymbol{W}_{q}-\boldsymbol{X}\boldsymbol{W}\|^{2},$ (1)
where $\mathcal{W}$ is an appropriate set of all feasible quantized weights. The matrix input $\boldsymbol{X}$ is the feature generated from the pre-trained model and does not depend on the quantized weights from the previous layer. We propose to solve the problem (1) by the coordinate descent algorithm. We regard all elements in the quantized weight and the scale factor as coordinates. In each step, we only solve a univariate optimization problem with respect to only one coordinate. By iteratively computing the quantized weights and corresponding scale factors, we finally find a proper quantized weight $\boldsymbol{W}_{q}$ that has a small quantization error and maintains good accuracy.
In this section, we will show our COMQ method for two quantization scenarios: per-layer quantization and per-channel quantization. All quantized weights with in the same layer share a common scale factor for the per-layer quantization, while different columns of the quantized weight use different scale factors for the per-channel quantization. Throughout the paper, we consider the $b$-bit asymmetric uniform quantization [4], which takes the bit-code set of $\mathbb{S}=\{z,z+1,\ldots,z+2^{b}-1\}$, with $z$ being the so-called zero point.
@@ 3.1 Per-layer Quantization
In this subsection, we present our COMQ method for $b$-bit per-layer quantization, where the whole quantized weight matrix $\boldsymbol{W}_{q}$ shares the scale factor $\delta$. In this setting, the quantized weight $\boldsymbol{W}_{q}$ can be decomposed as
$\boldsymbol{W}_{q}=\delta\cdot\boldsymbol{Q}\in\mathbb{R}^{m\times n},$
@@
where $\boldsymbol{Q}\in\mathbb{S}^{m\times n}$ is the integer bit-code matrix for $\boldsymbol{W}_{q}$. The optimization problem (1) reduces to
$\min_{\delta\in\mathbb{R},\boldsymbol{Q}\in\mathbb{S}^{m\times n}}\;\|\delta\boldsymbol{X}\boldsymbol{Q}-\boldsymbol{X}\boldsymbol{W}\|^{2}.$ (2)
@@
It holds that $\boldsymbol{X}\boldsymbol{Q}=(\boldsymbol{X}\boldsymbol{q}_{1},\ldots,\boldsymbol{X}\boldsymbol{q}_{n})$ and $\boldsymbol{X}\boldsymbol{W}=(\boldsymbol{X}\boldsymbol{w}_{1},\ldots,\boldsymbol{X}\boldsymbol{w}_{n})$, where $\boldsymbol{q}_{j}$ and $\boldsymbol{w}_{j}$ are the $j$-th column of $\boldsymbol{Q}$ and $\boldsymbol{W}$ respectively. Therefore, we can rewrite the problem (2) as
$\min_{\delta,\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{n}}\sum_{j=1}^{n}\|\delta\boldsymbol{X}\boldsymbol{q}_{j}-\boldsymbol{X}\boldsymbol{w}_{j}\|^{2}.$ (3)
@@
When fixing the scale factor $\delta$, the terms $\|\delta\boldsymbol{X}\boldsymbol{q}_{j}-\boldsymbol{X}\boldsymbol{w}_{j}\|^{2}$ for $j=1,\ldots,n$ are independent of each other. Hence, we can update the columns $\boldsymbol{q}_{1},\ldots,\boldsymbol{q}_{n}$ in parallel. We apply the coordinate descent method to solve the above problem (2). Specifically, we regard the scaling factor $\delta$ and all elements in $\boldsymbol{Q}$ as individual coordinates. In each iteration, we can update the bit-codes across all columns in the same row, i.e., row-wise update, before updating $\delta$:
$\{Q_{1,j}\}_{j\in[n]}\Rightarrow\{Q_{2,j}\}_{j\in[n]}\Rightarrow\cdots\Rightarrow\{Q_{m,j}\}_{j\in[n]}\Rightarrow\delta.$ (4)
@@
Initialization. At the beginning of our COMQ method, the scale factor $\delta^{0}$ should be properly initialized to capture the range of the weight matrix $\boldsymbol{W}$. Typically, the maximum value of $|\boldsymbol{W}|$ is the default selection in most quantization problems. However, this choice overlooks the impact of outliers in the weights on the overall quantization error. In order to smooth out these outliers, we consider the average infinity norm of weights across all columns $\boldsymbol{W}$. For the $b$ bits uniform quantization, we use the initialization $\delta^{0}=\frac{1}{2^{b-1}}\frac{\sum_{1\leq i\leq n}\|\boldsymbol{w}_{i}\|_{\infty}}{n}$. Then the matrix $\boldsymbol{Q}$ is initialized as $\boldsymbol{Q}^{0}=\frac{\boldsymbol{W}}{\delta^{0}}$. Note that the initialization $\boldsymbol{Q}^{0}$ above is not an actual bit-code matrix as $\boldsymbol{W}$ is float, but it will become feasible after the 1st iteration as will be shown below.
$\boldsymbol{Q}$-Update. Let $\boldsymbol{Q}^{k-1},\delta^{k-1}$ be the parameters after $k-1$ iterations. Without loss of generality, we focus on updating an arbitrary column of $\boldsymbol{Q}$, denoted by $\boldsymbol{q}$, where the column index $j$ is omitted for notational simplicity. Suppose the coordinates $q^{k}_{t}$ for $11\leq t\leq i-$ 1, have been updated in the $k$-th iteration. Note that $\boldsymbol{X}\boldsymbol{q}=\sum_{t=1}^{m}q_{t}\boldsymbol{x}_{t}$, with $\boldsymbol{x}_{t}$ being the $t$-th column of $\boldsymbol{X}$. Coordinate descent calls for solving the following problem to obtain $q^{k}_{i}$:
$q^{k}_{i}=$ $\;\arg\min_{q_{i}\in\mathbb{S}}\left\|\delta^{k-1}\left(q_{i}\boldsymbol{x}_{i}+\sum_{t=1}^{i-1}q_{t}^{k}\boldsymbol{x}_{t}+\sum_{t=i+1}^{m}q_{t}^{k-1}\boldsymbol{x}_{t}\right)-\boldsymbol{X}\boldsymbol{w}\right\|^{2}$
$=$ $\;\arg\min_{q_{i}\in\mathbb{S}}\left\|q_{i}\delta^{k-1}\boldsymbol{x}_{i}-\boldsymbol{s}_{i}^{k}\right\|^{2},$ (5)
@@
$q^{k}_{i}=\text{clip}\left(\left\lfloor\frac{\left<\boldsymbol{x}_{i},\boldsymbol{s}^{k}_{i}\right>}{\delta^{k-1}\|\boldsymbol{x}_{i}\|^{2}}\right\rceil,z,z+2^{b}-1\right).$
Hereby we define the quantization residuals at the $i$-th step (for the column):
$\boldsymbol{u}_{i}^{k}:=\sum_{t=1}^{i-1}(w_{t}-\delta^{k-1}q_{t}^{k})\boldsymbol{x}_{t}+\sum_{t=i}^{m}(w_{t}-\delta^{k-1}q_{t}^{k-1})\boldsymbol{x}_{t}=\boldsymbol{s}_{i}^{k}-\delta^{k-1}q_{i}^{k-1}.$
To efficiently implement the $\boldsymbol{Q}$-update as in (4), we take advantage of vectorized operations to update $\{Q_{i,j}\}_{j\in[n]}$ across all columns for each $i$. To this end, we denote by $\boldsymbol{w}_{i,:}\in\mathbb{R}^{n}$ and $\boldsymbol{q}_{i,:}\in\mathbb{R}^{n}$ the $i$-th row of $\boldsymbol{W}$ and $\boldsymbol{Q}$, respectively, and denote by $\boldsymbol{x}_{:,i}\in\mathbb{R}^{m}$ the $i$-th column of $\boldsymbol{X}$. Then the vectorized update of $\{Q_{i,j}^{k}\}_{j\in[n]}$ or $\boldsymbol{q}_{i,:}^{k}$ for the $k$-th iteration of COMQ proceeds as follows: with $\boldsymbol{U}_{0}^{k}=\boldsymbol{X}(\boldsymbol{W}-\delta^{k-1}\boldsymbol{Q}^{k-1})$, we iterate for $i=1,\dots,m$:
$\boldsymbol{U}_{i}^{k}=\boldsymbol{U}_{i-1}^{k}-\boldsymbol{x}_{:,i}\otimes(\boldsymbol{w}_{i,:}-\delta^{k-1}\boldsymbol{q}_{i,:}^{k-1})$
$\boldsymbol{q}^{k}_{i,:}=\text{clip}\left(\left\lfloor\frac{(\boldsymbol{U}^{k}_{i}+\boldsymbol{x}_{:,i}\otimes\boldsymbol{w}_{i,:})^{\top}\boldsymbol{x}_{:,i}}{\delta^{k-1}\boldsymbol{x}_{:,i}^{\top}\boldsymbol{x}_{:,i}}\right\rceil,z,z+2^{b}-1\right)$ (6)
$\boldsymbol{U}_{i}^{k}=\boldsymbol{U}_{i}^{k}+\boldsymbol{x}_{:,i}\otimes(\boldsymbol{w}_{i,:}-\delta^{k-1}\boldsymbol{q}_{i,:}^{k}).$
Here $\boldsymbol{U}_{i}^{k}$ maintains the quantization residuals across all columns.
$\delta$-Update. After obtaining the new bit-code matrix $\boldsymbol{Q}^{k}$, the scale factor can be updated by solving the following problem
$\min_{\delta}\|\delta\boldsymbol{X}\boldsymbol{Q}^{k}-\boldsymbol{X}\boldsymbol{W}\|^{2}.$
@@
$\delta^{k}=\frac{\left<\boldsymbol{X}\boldsymbol{Q}^{k},\boldsymbol{X}\boldsymbol{W}\right>}{\|\boldsymbol{X}\boldsymbol{Q}^{k}\|^{2}}.$ (7)
@@
Algorithm 1 COMQ for the per-layer quantization of one linear layer.
0:  Pre-trained weights $\boldsymbol{W}\in\mathbb{R}^{m\times n}$, feature matrix $\boldsymbol{X}$, and iteration number $K$.
1:  Initialize $\delta^{0}$ and $\boldsymbol{Q}^{0}$
@@
3:     for $i=1,\ldots,m$ do
4:        Update the coordinates $\{Q_{i,j}^{k}\}_{j\in[n]}$ or $\boldsymbol{q}^{k}_{i,:}$ as in (3.1)
@@
6:     Update the scaling factor $\delta^{k}$ as in (7)
@@
8:  Let $\boldsymbol{W}_{q}=\delta^{K}\boldsymbol{Q}^{K}$
8:  Quantized weight $\boldsymbol{W}_{q}$.